# Competition between transient oscillations and early stochasticity in exponentially growing populations

### v2

@@ IV Oscillation amplitude and phase vary across repeated identical experiments
In this section, we consider the case of $\epsilon<1$ and discuss the effects of fluctuations on the oscillations in single-population experiments. In Fig. 2 we see the trajectory of $\Lambda(t)$ for two populations with the same parameters and initial conditions. In all simulations, we observe transient oscillations of the form
@@
Although time scales $\Lambda_{\infty}$, $r$, and $\Omega$ are the same as for the ensemble’s oscillations, the amplitude $A$ and phase $\phi$ are now random variables that are set by the population’s early history, which is in close agreement with observations from Ref. [11]. In Fig. 3 (a) and (b), we show the spread in the instantaneous amplitude (see Appendix C for the definition) after 12 generations. Interestingly, the ensemble average amplitude $A_{\text{ens}}$ from Ref. [25] visibly underestimates the typical amplitude of a single population experiment.
Figure 2: Two trajectories of cell count growth rates $\Lambda(t)/\bar{\lambda}$ for two simulations with identical initial conditions showing different oscillation amplitudes at a long time. Simulation details: starting with one cell with mass $m(0)=\bar{m}$ and growth rate $\lambda_{0}=\bar{\lambda}$ with growth parameters $\tau_{\text{cor}}=5/\bar{\lambda}$, $\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}=0.003$ and division parameters $\text{CV}_{m_{d}}=0.12$, $\alpha=0.5$.
@@
In this section, we de-generalize our model a little to investigate how the amplitude prefactor $A$ in Eq. (12) is distributed and depends on the model’s parameters. We assume that the growth process $\lambda_{t}$ is Ornstein-Uhlenbeck with mean $\bar{\lambda}$, variance $\sigma_{\lambda}^{2}$, correlation time $\tau_{corr}$, and that the initial growth rate $\lambda_{0}$ is close to $\bar{\lambda}$. We also assume that the cell size at division admits some steady-state distribution with a coefficient of variation of $\text{CV}_{m_{b}}$. Based on the theory from Ref. [25], the ensemble average amplitude of a population starting from a single cell is approximately given by
@@
To summarize Eq. (13), two major factors affect the ensemble average amplitude. Firstly, growth rate correlations increase this amplitude, as they cause some delay in the time it takes for cells to fluctuate independently and de-synchronize their masses. Secondly, variability in division mass decreases the amplitude, as this causes a local spread in cell division timings among cells with synchronized mass, thereby slightly flattening oscillations in $\Lambda(t)$.
Figure 3: Instantaneous relative oscillation amplitudes in $\Lambda(t)$ for $4000$ simulations and their distribution. (a) and (b) represent the same set of simulations that started with a single cell, while in (c) and (d), the population started with three cells. In (a) and (c), the black line is the root mean square $\sqrt{\langle A^{2}\rangle}$, and the green line is its asymptotic prediction $\bar{A}e^{-rt}$, where $\bar{A}$ is given by Eq. (15). The brown solid line is the ensemble average’s amplitude $|\langle Ae^{i\phi}\rangle|$, and the red dashed line is its asymptotic prediction $A_{\text{ens}}e^{-rt}$, where $A_{\text{ens}}$ is given by Eq. (13). In (b) and (d) we show the distribution of the amplitude pre-factor $A$ based on their values at $t=12\tau_{\text{div}}$. Simulations in (c) and (d) start with three cells with masses $m(0)/\bar{m}=1,1.19,1.56$, chosen such that initial cell mass synchronicity is $\Psi(0)=0$.
@@
In Appendix C we show that this typical amplitude is closely approximated by
@@
In Fig. 3 (a) and (b) we see how the typical amplitude pre-factor $\bar{A}$ more closely resembles the amplitude in typical experiments than $A_{\text{ens}}$. The right-most factor in Eq. (15) is strictly greater than one, and it quantifies the net increase in typical amplitude pre-factor with respect to the ensemble average due to stochasticity in the early population. In Fig. 1 (a), we plot $\bar{A}$ for various values of $\epsilon$ and $\tau_{\text{cor}}$. Note that $\bar{A}$ diverges at the oscillation-fluctuation phase transition as $\epsilon$ goes to $1$. This does not mean that we observe arbitrarily large oscillations, since an increase in $\epsilon$ also means that oscillations decay faster. The instantaneous oscillation amplitude at a fixed point in time therefore still goes to a constant as $\epsilon$ goes to $1$.
@@
We have seen in Eq. (15) that the typical amplitude $\bar{A}$ diverges as we approach $\epsilon=1$ (see also Fig. 4 (a)). We can understand this by interpreting $A$ in Eq. (12) as the Fourier amplitude $e^{rt}(\Lambda(t)-\Lambda_{\infty})$ at frequency $\Omega$. As we approach the transition point, the contribution of the neglected fluctuations in Eq. (12) to this Fourier mode increases and diverges after the transition point. This is because the neglected noise decays as $e^{-\Lambda_{\infty}/2}$ and when multiplied by $e^{rt}$ it becomes a growing function of time for $r>\Lambda_{\infty}/2$ (i.e. $\epsilon>1$). On the other hand, the ensemble amplitude $A_{ens}$ is a measure of this Fourier amplitude after the fluctuations in $N(t)$ are averaged out. With this in mind, we can interpret the ratio $A_{ens}/\bar{A}$ as a measure of how much of this Fourier mode is dominated by the deterministic oscillations. It is a measure of the order in this Fourier mode, sharply decays to zero at $\epsilon=1$, and stays zero after the transition point. It plays the same role that the order parameter plays in equilibrium phase transitions.
Moreover, the dependence of $A_{ens}$ and $\bar{A}$ on all other model parameters beside $\epsilon$ cancel out when dividing Eq. (13) by Eq. (15) giving the simple universal dependence of the order parameter on $\epsilon$ given by
@@
Figure 4 shows the dependence of $A_{ens}$ and $\bar{A}$ and their ratio on $\epsilon$ and $\tau_{cor}$ (theory and simulation), showing the ratio collapses all the curves into the single curve given by Eq. (16).
Figure 4: Analytically calculated (solid curves) dependence of the typical amplitude $\bar{A}$, ensemble amplitude $A_{ens}$, and their ratio on the relative decay rate $\epsilon$ for various values of correlation times $\tau_{cor}$ compared to simulation results (markers). (a) The typical amplitude $\bar{A}$ has a strong dependence on $\epsilon$ (given in Eq. (15)) that diverges at $\epsilon=1$, and it decreases with $\tau_{cor}$. (b) The ensemble amplitude $A_{ens}$ has a relatively moderate dependence on $\epsilon$ and $\tau_{cor}$ (given in Eq. (13)). (c) When divided by each other, the $\epsilon$-dependence collapses on a single curve that is zero after the transition point $\epsilon=1$ and is given by Eq. (16) before the transition. In that sense, the ratio $A_{ens}/\bar{A}$ is analogous to the order parameter in equilibrium phase transitions and is a measure of how deterministic the oscillations look. Note that numerically probing the vicinity of the transition point is more challenging as the required simulation time diverges at the transition point, and the number of cells grows exponentially with time. Simulation details: the pre-factors were estimated as $A\approx 2|\Psi(t)|e^{rt}$ (see Appendix C), for large simulation time $t$. $\text{CV}_{m_{d}}$ is set to zero. For $\epsilon=0.1,0.3,0.5,0.7$, we used $1000$ independent simulations and a simulation time of $t=17\tau_{\text{div}}$. For $\epsilon=0.9$, we used $200$ independent simulations and a simulation time of $t=21\tau_{\text{div}}$.
@@ V Oscillation amplitude in a population starting from multiple cells
The absolute value $|\Psi(t)|$ ranges between 00 and $1$ and measures the degree to which cell masses are synchronized. When $|\Psi(t)|=1$, all cells must have the same mass. In the special case of one cell, this always holds, hence Eq. (17) reduces to Eq. (13). When all cells are out of phase, we get $\Psi(t)=0$. In this case, there is full destructive interference in the first-order ensemble average oscillations of the future trajectory of $\Lambda_{\text{ens}}(t)$. Naturally, we have $\Psi(t)=0$ when cell masses are distributed according to the steady-state population mass distribution [25]. Whenever we talk about the instantaneous amplitude, we actually mean a re-scaled $|\Psi(t)|$. The general form of the typical amplitude is
@@
The larger the initial population, the smaller $J(0)$ and the smaller the effect of early population stochasticity. The strength of this effect also monotonically increases with growth rate variability. Interestingly, $J(0)$ is always a positive finite value, so $\bar{A}$ is strictly positive, even when the initial population is completely out of phase. This is demonstrated in Fig. 3 (c) and (d), where we provide the trajectories of instantaneous amplitudes as well as their distributions for a population starting from three out-of-phase cells. Note how in Fig. 3 (c), the small population immediately spontaneously synchronizes its cell masses, which then leads to sustained oscillations in the transient regime.
@@ VI How long do the oscillations last?
Let us estimate what this amounts to for some experimentally obtained parameters, given a population that starts from a single cell with mass $\bar{m}$ at $t=0$. In Ref. [27], single lineage data is collected for E. coli in various growth media. We are interested in the mean and coefficient of variation of the elongation rate and cell length at division, as well as the elongation rate’s mother-daughter correlations. We assume cell length is proportional to cell mass and that the reported elongation rates are effective cell-cycle averages of some underlying Ornstein-Uhlenbeck growth rate process. Using theory from Ref. [25], we can deduce the set of parameters $\bar{\lambda}$, $\sigma_{\lambda}$, and $\tau_{\text{cor}}$ of the supposed underlying growth rate process, see Appendix D for the relationships used. In Table 1 we report these values, derived from the experiments obtained over the seven different growth media. We also calculate the oscillation decay rate $r$, order parameter $\epsilon$, and typical amplitude $\bar{A}$. At last, we report the typical number of generations a population needs for the relative amplitude to fall below a cutoff of $0.1$. Note the large discrepancy ranging between $12$ and $34$ generations in this cutoff. This corresponds to a million-fold increase in population number. The growth medium thus has a large effect on whether the oscillations are observable or not.
@@
Table 1: Experimentally measured values of the model parameters for various growth conditions for E. coli obtained from Ref. [27] (see Appendix D for details) together with analytically calculated values for oscillation decay rate $r$, relative decay rate $\epsilon$, typical amplitude pre-factor $\bar{A}$, and number of generation before the amplitude falls below $10\%$ of average growth rate. All seven growth conditions are below the transition points with "glucose + 12 amino acids" having the longest-lasting oscillations with amplitude larger than $10\%$, lasting for over 34 generations. See Ref. [27] for the details of the growth conditions.
@@ VII Discussions and future directions
In the 1950s, well before the existence of single-cell technology, microbiologists resorted to developing experimental methods to synchronize populations of growing bacteria to be able to probe single cells [32, 33]. A celebrated example is what was widely known as the “baby machine” that screened cells of the same size to produce a synchronized population [34]. This method was famously used to discover multi-fork DNA replication in E. coli [35]. These experiments inspired many early attempts to theoretically model the dynamics of synchronized growing populations [21, 36, 37, 20]. These attempts use McKendrick-von Foerster formalism to analyze models akin to Bellman-Harris [38], where cells grow for a period of time independently drawn from a division time distribution and then divide. Such models are sometimes referred to as timer models or independent generation time models. These models predict that variation in generation times of cells (about $20\%$ variability [27]) desynchronizes the population (note that $20\%$ is well above the $13\%$ threshold in Eq. (1), so no one would have expected to get a synchronized population just by starting it from a single cell).
@@
With the recent invention of single-cell Microfluidics technology such as the "mother machine" [39] and subsequent theoretical developments in cell size regulation [31], we now understand that the aforementioned theoretical models fail to capture the oscillation dynamics quantitatively. This is because cell-size regulation drastically slows down the desynchronization [1]. There are two major sources of variability in the generation time of a cell: the variability in the growth rate, and the variability in the size at which the cells divide. The decay rate of the oscillations is set by the smaller source of noise, the noise in the growth rate of the cells. This is while, any long-term effect of the larger source of noise, the noise in the division size, on the population dynamics, is canceled by cell-size regulation. This comes with the realization that under conditions where the growth rate variability is sufficiently low, these oscillations can last longer than 30 generations. This is the time it takes for a $1$ ml culture, starting with a single cell, to saturate. In other words, all one needs to synchronize a large population is to start it from a single cell. This provides a way to probe hard-to-measure single-cell statistics by performing simple population measurements.
@@
There have been very few experimental works on measurements of oscillations in population starting from single cells, and they only capture the very first few generations on gel plates using simple microscopy [13, 18]. Our theoretical predictions suggest that for many growth conditions, these oscillations can be very easily measured in liquid culture at the late stage when the population is growing deterministically. For this reason, we have written this paper with our experimental colleagues in mind to provide the necessary tools that help decide when these oscillations can be observed.
@@
It is important to realize that these oscillations are hidden from typical optical density measurements which estimate the total mass and not directly the cell count. The period of these oscillations is about one generation time, which means one needs multiple repeated measurements of instantaneous population growth rate per cell cycle (which could be as short as 23 min for E. coli). The issue of time resolution can be alleviated by performing the experiment at a lower temperature, where all the processes in cells slow down almost proportionally. The population growth rate can be estimated by cell counting measurements performed twice on small samples a few minutes apart. We expect any well-designed experiment to be able to estimate the instantaneous population growth rate with uncertainty well below $10\%$. For this reason, we have estimated in Table 1 the number of generations it takes before the amplitude of the oscillations falls below $10\%$ of the mean growth rate. For certain growth conditions (glucose + 12 amino acids, see Ref. [27] for the description of the growth media), the oscillations are predicted to maintain amplitude above $10\%$ of the mean growth rate for $34$ generations (this corresponds to longer than the time it takes a $10$ ml culture to saturate).
@@
Here, we have shown that when the relative decay rate $\epsilon$ goes above 1, these oscillations are not observable. This threshold corresponds to roughly $13\%$ variability in growth rate, which is within the range $6-20\%$ variability reported in the literature [27, 40, 41, 42]. However, careful estimations of the parameters of the model for all seven growth conditions from Ref. [27] put E. coli below this threshold (see Table 1). We do not know if there is a biological significance for cells to stay below this transition point, and we have no reason to believe this is always the case. Moreover, due to the finite uncertainty in any measurement of cell counts, we find it unlikely for these oscillations to be observed in single experiments when $\epsilon$ is close to one which is the case for some of the growth conditions.
@@
In this work, we have tried to use a realistic model that keeps track of noise in both growth and division, is compatible with models of cell size regulation, and includes correlations in growth rates. One can always write more complicated models of growth and division keeping track of more degrees of freedom, but complicated models come with more parameters. We believe our model captures all the dominant sources of noise and correlations without involving parameters that are unknown or hard to measure. One source of noise and correlation that is being ignored in our model is any effect of asymmetry in cell division. E. coli is extremely good at dividing symmetrically; the asymmetry in cell division of E. coli is measured to have the relative variability of order $2\%$ under certain conditions [30], which would certainly have a negligible effect on any calculations performed in this paper. However, one could expect growth conditions or other species of bacteria in which there is a significant source of noise in the position of the division plane. The asymmetry in cell division could also manifest itself in an asymmetry in the division of the proteins in ribosomes among the two daughter cells which would lead to discontinuity in the growth process and subtle correlations between the growth rate and cell size [43, 44]. Quantifying the effect of such asymmetries would be the subject of future work.

### v3

@@ IV Oscillation amplitude and phase vary across repeated identical experiments
In this section, we consider the case of $\epsilon<1$ and discuss the effects of fluctuations on the oscillations in single-population experiments. Figure 2 (a) and (b) show the trajectories of $\Lambda(t)$ for two populations with the same parameters and initial conditions. In all simulations, we observe transient oscillations of the form
@@
Although time scales $\Lambda_{\infty}$, $r$, and $\Omega$ are the same as for the ensemble’s oscillations, the amplitude $A$ and phase $\phi$ are now random variables that are set by the population’s early history, which is in close agreement with observations from Ref. Gavagnin et al. (2021). In Fig. 3 (a) and (b), we show the spread in the instantaneous amplitude (see Appendix D for the definition) after 12 generations. Interestingly, the ensemble average amplitude $A_{\text{ens}}$ from Ref. Hein and Jafarpour (2022) visibly underestimates the typical amplitude of a single population experiment.
Figure 2: (a) and (b) show two trajectories of cell count growth rates $\Lambda(t)/\bar{\lambda}$ for two simulations with identical parameters and initial conditions. Note how the oscillations have different amplitudes despite both simulations having identical initial conditions. Simulation details: starting with one cell with mass $m(0)=\bar{m}=1$ and growth rate $\lambda_{0}=\bar{\lambda}=\ln(2)$ with growth parameters $\tau_{\text{cor}}=5/\bar{\lambda}$, $\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}=0.003$ and division parameters $\text{CV}_{m_{d}}=0.12$, $\alpha=0.5$. This corresponds to a decay ratio of $\epsilon=0.5$.
@@
In this section, we de-generalize our model a little to investigate how the amplitude prefactor $A$ in Eq. (12) is distributed and depends on the model’s parameters. We assume that the growth process $\lambda_{t}$ is Ornstein-Uhlenbeck with mean $\bar{\lambda}$, variance $\sigma_{\lambda}^{2}$, correlation time $\tau_{corr}$, and that the initial growth rate $\lambda_{0}$ is close to $\bar{\lambda}$. We also assume that the cell size at division admits some steady-state distribution with a coefficient of variation of $\text{CV}_{m_{b}}$. Based on the theory from Ref. Hein and Jafarpour (2022), the ensemble average amplitude from Eq. (7) of a population starting from a single cell is approximately given by
@@
To summarize Eq. (13), two major factors affect the ensemble average amplitude. Firstly, growth rate correlations increase this amplitude, as they cause some delay in the time it takes for cells to fluctuate independently and de-synchronize their masses. Secondly, variability in division mass decreases the amplitude, as this causes a local spread in cell division timings among cells with synchronized mass, thereby slightly flattening oscillations in $\Lambda(t)$. In Fig. 4 (a), we plot $A_{\text{ens}}$ for various values of $\tau_{\text{cor}}$ and $\epsilon$ (using Eq. (11) to determine $\sigma_{\lambda}^{2}$) and no cell division variability $\text{CV}_{m_{b}}=0$. We see the dependence of $A_{\text{ens}}$ on the growth rate parameters is fairly limited in the relevant regime.
Figure 3: Instantaneous relative oscillation amplitudes in $\Lambda(t)$ for $4000$ simulations and their distribution. (a) and (b) represent the same set of simulations that started with a single cell, while in (c) and (d), the population started with three cells. The instantaneous amplitude is given by $A_{\text{inst}}(t)=e^{-41\text{CV}_{m_{d}}^{2}}|\Psi(t)|$, with $\Psi(t)$ given in Eq. (18). In (a) and (c), the black line is the root mean square $\sqrt{\langle A_{\text{inst}}(t)^{2}\rangle}$, and the green line is its asymptotic prediction $\bar{A}e^{-rt}$, where $\bar{A}$ is given by Eq. (15). The brown solid line is the ensemble average’s amplitude, and the red dashed line is its asymptotic prediction $A_{\text{ens}}e^{-rt}$, where $A_{\text{ens}}$ is given by Eq. (13). In (b) and (d) we show the distribution of the amplitude pre-factor $A_{\text{inst}}(t_{max})e^{rt_{max}}$ based on their values at $t_{max}=12\tau_{\text{div}}$. Simulations in (c) and (d) start with three cells with masses $m(0)/\bar{m}=1,1.19,1.56$, chosen such that initial cell mass synchronicity is $\Psi(0)=0$. The remaining simulation details are identical to the simulations used to generate Fig. 1 (c) (d) and Fig. 2.
@@
In Appendix D we show that this typical amplitude is closely approximated by
@@
In Fig. 3 (a) and (b) we see how the typical amplitude pre-factor $\bar{A}$ more closely resembles the amplitude in typical experiments than $A_{\text{ens}}$. The right-most factor in Eq. (15) is strictly greater than one, and it quantifies the net increase in typical amplitude pre-factor with respect to the ensemble average due to stochasticity in the early population. In Fig. 4 (b), we plot $\bar{A}$ for various values of $\epsilon$ and $\tau_{\text{cor}}$, showing that $\bar{A}$ diverges at the oscillation-fluctuation phase transition as $\epsilon$ goes to $1$. Note that this does not mean that we observe arbitrarily large oscillations, since an increase in $\epsilon$ also means that oscillations decay faster. The instantaneous oscillation amplitude at a fixed point in time therefore still goes to a constant as $\epsilon$ goes to $1$.
@@
We have seen in Eq. (15) that the typical amplitude $\bar{A}$ diverges as we approach $\epsilon=1$ (see also Fig. 4 (b)). We can understand this by interpreting $A$ in Eq. (12) as the Fourier amplitude $e^{rt}(\Lambda(t)-\Lambda_{\infty})$ at frequency $\Omega$. As we approach the transition point, the contribution of the neglected fluctuations in Eq. (12) to this Fourier mode increases and diverges after the transition point. This is because the neglected noise decays as $e^{-\Lambda_{\infty}/2}$ and when multiplied by $e^{rt}$ it becomes a growing function of time for $r>\Lambda_{\infty}/2$ (i.e. $\epsilon>1$). On the other hand, the ensemble amplitude $A_{ens}$ is a measure of this Fourier amplitude after the fluctuations in $N(t)$ are averaged out. With this in mind, we can interpret the ratio $A_{ens}/\bar{A}$ as a measure of how much of this Fourier mode is dominated by the deterministic oscillations. It is a measure of the order in this Fourier mode, sharply decays to zero at $\epsilon=1$, and stays zero after the transition point. It plays the same role that the order parameter plays in equilibrium phase transitions.
Figure 4 shows the dependence of $A_{ens}$ and $\bar{A}$ and their ratio on $\epsilon$ and $\tau_{cor}$ (theory and simulation). Strikingly, the dependence of the “order parameter”$A_{ens}/\bar{A}$ on $\epsilon$ collapses for various model parameters into the same curve. The dependence of $A_{ens}$ and $\bar{A}$ on all other model parameters beside $\epsilon$ cancel out when dividing Eq. (13) by Eq. (15) giving the simple universal dependence of the order parameter on $\epsilon$ given by
@@
This universal shape of the “order parameter”independent of model parameter exhibiting a sharp transition is analogous to the equilibrium phase transition. Similar to the equilibrium case, this sharp transition can only be observed in simulations at infinite system size (which corresponds to the large time limit in our growing system). Here, we have a nonequilibrium dynamics with a growing system size, for which (to the best of our knowledge) no theory of phase transition exists. All the comparisons with the equilibrium phase transitions are simple analogies. Some readers may prefer the language of abrupt nonlinear response to describe this behavior.
Figure 4: Analytically calculated (solid curves) dependence of the typical amplitude $\bar{A}$, ensemble amplitude $A_{ens}$, and their ratio on the relative decay rate $\epsilon$ for various values of correlation times $\tau_{cor}$ compared to simulation results (markers). (a) The ensemble amplitude $A_{ens}$ has a relatively moderate dependence on $\epsilon$ and $\tau_{cor}$ (given in Eq. (13)). (b) The typical amplitude $\bar{A}$ has a strong dependence on $\epsilon$ (given in Eq. (15)) that diverges at $\epsilon=1$, and it decreases with $\tau_{cor}$. (c) When divided by each other, the $\epsilon$-dependence collapses on a single curve that is zero after the transition point $\epsilon=1$ and is given by Eq. (16) before the transition. In that sense, the ratio $A_{ens}/\bar{A}$ is analogous to the order parameter in equilibrium phase transitions and is a measure of how deterministic the oscillations look. Note that numerically probing the vicinity of the transition point is more challenging as the required simulation time diverges at the transition point, and the number of cells grows exponentially with time. Simulation details: the pre-factors were estimated as $A\approx 2|\Psi(t)|e^{rt}$ (see Appendix D), for large simulation time $t$. $\text{CV}_{m_{d}}$ is set to zero. For $\epsilon=0.1,0.3,0.5,0.7$, we used $1000$ independent simulations and a simulation time of $t=17\tau_{\text{div}}$. For $\epsilon=0.9$, we used $200$ independent simulations and a simulation time of $t=24\tau_{\text{div}}$.
@@ V Oscillation amplitude in a population starting from multiple cells
The absolute value $|\Psi(t)|$ ranges between 00 and $1$ and measures the degree to which cell masses are synchronized. When $|\Psi(t)|=1$, all cells must have the same mass. In the special case of one cell, this always holds, hence Eq. (17) reduces to Eq. (13). When all cells are out of phase, we get $\Psi(t)=0$. In this case, there is full destructive interference in the first-order ensemble average oscillations of the future trajectory of $\Lambda_{\text{ens}}(t)$. Naturally, we have $\Psi(t)=0$ when cell masses are distributed according to the steady-state population mass distribution Hein and Jafarpour (2022). Whenever we talk about the instantaneous amplitude, we actually mean a re-scaled $|\Psi(t)|$. The general form of the typical amplitude is
@@
The larger the initial population, the smaller $J(0)$ and the smaller the effect of early population stochasticity. The strength of this effect also monotonically increases with growth rate variability. Interestingly, $J(0)$ is always a positive finite value, so $\bar{A}$ is strictly positive, even when the initial population is completely out of phase. This is demonstrated in Fig. 3 (c) and (d), where we provide the trajectories of instantaneous amplitudes as well as their distributions for a population starting from three out-of-phase cells. Note how in Fig. 3 (c), the small population immediately spontaneously synchronizes its cell masses, which then leads to sustained oscillations in the transient regime. When there is little to no synchronization in the initial population, we find that the asymptotic amplitude pre-factor $A$ of the later population follows a $\chi_{2}$ distribution, as discussed in Appendix D.1. In Fig. 3 (d) we see that this distribution matches the simulated oscillation pre-factors very well. The instantaneous amplitude is defined as $A_{inst}(t)=e^{-41\text{CV}_{m_{d}}}|\Psi(t)|$ and it is the best predictor for the amplitude of first order oscillations (see Appendix C.2 for a derivation and discussion).
@@ VI How long do the oscillations last?
Let us estimate what this amounts to for some experimentally obtained parameters, given a population that starts from a single cell with mass $\bar{m}$ at $t=0$. In Ref. Taheri-Araghi et al. (2015), single lineage data is collected for E. coli in various growth media. We are interested in the mean and coefficient of variation of the elongation rate and cell length at division, as well as the elongation rate’s mother-daughter correlations. We assume cell length is proportional to cell mass and that the reported elongation rates are effective cell-cycle averages of some underlying Ornstein-Uhlenbeck growth rate process. Using theory from Ref. Hein and Jafarpour (2022), we can deduce the set of parameters $\bar{\lambda}$, $\sigma_{\lambda}$, and $\tau_{\text{cor}}$ of the supposed underlying growth rate process, see Appendix E for the relationships used. In Table 1 we report these values, derived from the experiments obtained over the seven different growth media. We also calculate the oscillation decay rate $r$, relative decay rate $\epsilon$, and typical amplitude $\bar{A}$. At last, we report the typical number of generations a population needs for the relative amplitude to fall below a cutoff of $0.1$. Note the large discrepancy ranging between $12$ and $34$ generations in this cutoff. This corresponds to a million-fold increase in population number. The growth medium thus has a large effect on whether the oscillations are observable or not.
@@
Table 1: Experimentally measured values of the model parameters for various growth conditions for E. coli obtained from Ref. Taheri-Araghi et al. (2015) (see Appendix E for details) together with analytically calculated values for oscillation decay rate $r$, relative decay rate $\epsilon$, typical amplitude pre-factor $\bar{A}$, and number of generation before the amplitude falls below $10\%$ of average growth rate. All seven growth conditions are below the transition points with "glucose + 12 amino acids" having the longest-lasting oscillations with amplitude larger than $10\%$, lasting for over 34 generations. See Ref. Taheri-Araghi et al. (2015) for the details of the growth conditions.
@@ VII Discussions and future directions
In the 1950s, well before the existence of single-cell technology, microbiologists resorted to developing experimental methods to synchronize populations of growing bacteria to be able to probe single cells Lark and Maaløe (1954); Jun et al. (2018). A celebrated example is what was widely known as the “baby machine” that screened cells of the same size to produce a synchronized population Helmstetter and Cummings (1963). This method was famously used to discover multi-fork DNA replication in E. coli Cooper and Helmstetter (1968). These experiments inspired many early attempts to theoretically model the dynamics of synchronized growing populations Hirsch and Engelberg (1966); Bronk et al. (1968); Brown (1940); Kendall (1948). These attempts use McKendrick-von Foerster formalism to analyze models akin to Bellman-Harris Bellman and Harris (1952), where cells grow for a period of time independently drawn from a division time distribution and then divide. Such models are sometimes referred to as timer models or independent generation time models. These models predict that variation in generation times of cells (about $20\%$ variability Taheri-Araghi et al. (2015)) desynchronizes the population (note that $20\%$ is well above the $13\%$ threshold in Eq. (1), so no one would have expected to get a synchronized population just by starting it from a single cell).
@@
With the recent invention of single-cell Microfluidics technology such as the "mother machine" Wang et al. (2010) and subsequent theoretical developments in cell size regulation Amir (2014), we now understand that the aforementioned theoretical models fail to capture the oscillation dynamics quantitatively. This is because cell-size regulation drastically slows down the desynchronization Jafarpour (2019). There are two major sources of variability in the generation time of a cell: the variability in the growth rate, and the variability in the size at which the cells divide. The decay rate of the oscillations is set by the smaller source of noise, the noise in the growth rate of the cells. This is while, any long-term effect of the larger source of noise, the noise in the division size, on the population dynamics, is canceled by cell-size regulation. This comes with the realization that under conditions where the growth rate variability is sufficiently low, these oscillations can last longer than 30 generations. This is the time it takes for a $1$ ml culture, starting with a single cell, to saturate. In other words, all one needs to synchronize a large population is to start it from a single cell. This provides a way to probe hard-to-measure single-cell statistics by performing simple population measurements.
@@
There have been very few experimental works on measurements of oscillations in population starting from single cells, and they only capture the very first few generations on gel plates using simple microscopy Chiorino et al. (2001); Vittadello et al. (2019). Our theoretical predictions suggest that for many growth conditions, these oscillations can be very easily measured in liquid culture at the late stage when the population is growing deterministically. For this reason, we have written this paper with our experimental colleagues in mind to provide the necessary tools to help decide when these oscillations can be observed.
@@
It is important to realize that these oscillations are hidden from typical optical density measurements which estimate the total mass and not directly the cell count. The period of these oscillations is about one generation time, which means one needs multiple repeated measurements of instantaneous population growth rate per cell cycle (which could be as short as 23 min for E. coli). The issue of time resolution can be alleviated by performing the experiment at a lower temperature, where all the processes in cells slow down almost proportionally. The population growth rate can be estimated by cell counting measurements performed twice on small samples a few minutes apart. We expect any well-designed experiment to be able to estimate the instantaneous population growth rate with uncertainty well below $10\%$. For this reason, we have estimated in Table 1 the number of generations it takes before the amplitude of the oscillations falls below $10\%$ of the mean growth rate. For certain growth conditions (glucose + 12 amino acids, see Ref. Taheri-Araghi et al. (2015) for the description of the growth media), the oscillations are predicted to maintain amplitude above $10\%$ of the mean growth rate for $34$ generations (this corresponds to longer than the time it takes a $10$ ml culture to saturate).
@@
Here, we have shown that when the relative decay rate $\epsilon$ goes above 1, these oscillations are not observable. This threshold corresponds to roughly $13\%$ variability in growth rate, which is within the range $6-20\%$ variability reported in the literature Taheri-Araghi et al. (2015); Cermak et al. (2016); Kennard et al. (2016); Wallden et al. (2016). However, careful estimations of the parameters of the model for all seven growth conditions from Ref. Taheri-Araghi et al. (2015) put E. coli below this threshold (see Table 1). We do not know if there is a biological significance for cells to stay below this transition point, and we have no reason to believe this is always the case. Moreover, due to the finite uncertainty in any measurement of cell counts, we find it unlikely for these oscillations to be observed in single experiments when $\epsilon$ is close to one which is the case for some of the growth conditions.
@@
In this work, we have tried to use a realistic model that keeps track of noise in both growth and division, is compatible with models of cell size regulation, and includes correlations in growth rates. One can always write more complicated models of growth and division keeping track of more degrees of freedom, but complicated models come with more parameters. We believe our model captures all the dominant sources of noise and correlations without involving parameters that are unknown or hard to measure. One source of noise and correlation that is being ignored in our model is any effect of asymmetry in cell division. E. coli is extremely good at dividing symmetrically; the asymmetry in cell division of E. coli is measured to have the relative variability of order $2\%$ under certain conditions Guberman et al. (2008), which would certainly have a negligible effect on any calculations performed in this paper. However, one could expect growth conditions or other species of bacteria in which there is a significant source of noise in the position of the division plane. The asymmetry in cell division could also manifest itself in an asymmetry in the division of the proteins in ribosomes among the two daughter cells which would lead to discontinuity in the growth process and subtle correlations between the growth rate and cell size Harsh et al. (2021); Nordholt et al. (2020). Quantifying the effect of such asymmetries would be the subject of future work.